# Do LLMs Have Distinct and Consistent Personality? TRAIT: Personality Testset designed for LLMs with Psychometrics

### v1

@@ G.1 Refusal Rate
Test Prompt Type Option Choice Sensitivity Binary Level Sensitivity
Type 1 Type 2 Type 3 Average (Std) Type 1 Type 2 Type 3 Average (Std)
BIG-5 GPT-4 27.3 18.2 27.3 24.3 (4.29) 11.4 9.1 6.8 9.1 (1.9)
GPT-3.5 84.1 40.9 97.7 74.2 (24.2) 84.1 18.2 93.2 65.2 (33.4)
Llama3-8B-instruct 77.3 81.8 93.2 84.1 (6.7) 72.7 43.2 93.2 69.7 (20.5)
Llama2-7B-chat 100.0 100.0 93.2 97.7 (3.2) 100.0 100.0 93.2 97.7 (3.2)
Mistral-7B-instruct 95.5 97.7 27.3 73.5 (32.7) 52.3 79.5 22.7 51.5 (23.2)
Mistral-7B-sft 70.5 100.0 100.0 90.2 (13.9) 56.8 100.0 100.0 85.6 (20.4)
Tulu2-7B-DPO 68.2 88.6 100.0 85.6 (13.2) 63.6 75.0 100.0 79.5 (15.2)
SD-3 GPT-4 59.3 0.0 33.3 30.9 (24.3) 25.9 0.0 3.7 9.9 (11.4)
GPT-3.5 66.7 51.9 92.6 70.4 (16.8) 66.7 37.0 92.6 65.4 (22.7)
Llama3-8B-instruct 96.3 51.9 77.8 75.3 (18.2) 74.1 25.9 77.8 59.3 (23.6)
Llama2-7B-chat 100.0 100.0 92.6 97.5 (3.5) 100.0 100.0 92.6 97.5 (3.5)
Mistral-7B-instruct 96.3 96.3 25.9 72.8 (33.2) 14.8 51.9 22.2 29.6 (16.0)
Mistral-7B-sft 85.2 100.0 100.0 95.1 (7.0) 59.3 100.0 100.0 86.4 (19.2)
Tulu2-7B-DPO 88.9 92.6 100.0 93.8 (4.6) 81.5 77.8 100.0 86.4 (9.7)
IPIP-NEO-PI GPT-4 39.7 9.0 37.7 28.8 (14.0) 26.3 2.3 10.3 13 (10.0)
GPT-3.5 72.3 47.0 94.0 71.1 (19.2) 70.3 28.7 89.3 62.8 (25.3)
Llama3-8B-instruct 93.3 80.7 83.3 85.8 (5.4) 80.7 39.0 83.3 67.7 (20.3)
Llama2-7B-chat 100.0 100.0 99.7 99.9 (0.1) 99.3 100.0 99.7 99.7 (0.3)
Mistral-7B-instruct 97.7 97.0 31.0 75.2 (31.3) 26.0 52.3 31.0 36.4 (11.4)
Mistral-7B-sft 77.7 100.0 100.0 92.6 (10.5) 43.0 100.0 100.0 81.0 (26.87)
Tulu2-7B-DPO 76.7 92.0 100.0 89.6 (9.7) 68.7 74.7 99.7 81.0 (13.4)
Anthropic-Eval GPT-4 0.8 0.0 0.0 0.3 (0.4) 0.8 0.0 0.0 0.27 (0.4)
GPT-3.5 7.5 6.4 13.5 9.1 (3.1) 7.5 6.4 13.5 9.1 (3.1)
Llama3-8B-instruct 7.7 15.6 56.4 26.6 (21.3) 7.7 15.6 56.4 26.6 (21.3)
Llama2-7B-chat 65.0 46.0 100.0 70.3 (22.4) 65.0 46.0 100.0 70.3 (22.4)
Mistral-7B-instruct 26.1 36.5 86.7 49.8 (26.5) 26.1 36.5 86.7 49.8 (26.5)
Mistral-7B-sft 41.3 48.6 100.0 63.3 (26.1) 41.3 48.6 100.0 63.3 (26.1)
Tulu2-7B-DPO 57.3 60.0 75.0 64.1 (7.8) 57.3 60.0 75.0 64.1 (7.8)
TRAIT GPT-4 29.8 26.9 27.0 27.9 (1.3) 10.7 6.7 9.1 8.8 (1.6)
GPT-3.5 39.7 20.1 38.8 32.9 (9.0) 23.5 8.0 21.7 17.7 (6.9)
Llama3-8B-instruct 76.1 61.4 98.6 78.7 (15.3) 45.5 34.6 96.1 58.7 (26.8)
Llama2-7B-chat 43.7 33.7 40.9 39.4 (4.2) 34.0 23.1 25.6 27.6 (4.7)
Mistral-7B-instruct 39.4 33.3 59.1 43.9 (11.0) 23.7 19.2 48.3 30.4 (12.8)
Mistral-7B-sft 51.8 46.4 94.0 64.1 (21.3) 28.9 23.7 69.6 40.7 (20.5)
Tulu2-7B-DPO 43.3 45.1 55.5 48 (5.4) 19.6 12.9 31.4 21.3 (7.7)
@@ I.1 Labelers
In Tables 20, 21, and 22, we display the qualitative responses from GPT-4. These responses are from different questionnaires, starting with the same personality descriptions.
@@ 4. Why not just chill and watch a movie together when you have time?
Figure 16: Labeling interface which is shown to psychological professionals. We utilize label-studio888label-studio in making a simple interface.

### v2

@@ G.1 Refusal Rate
Test Model Option Choice Sensitivity
Type 1 Type 2 Type 3 Average
BFI GPT-3.5 0.0 20.5 79.5 33.3
Mistral-7B-instruct 47.7 72.7 56.8 59.1
Mistral-7B-sft 31.8 100.0 100.0 77.3
Llama3-8B-instruct 97.7 45.5 22.7 55.3
Tulu2-7B 65.9 100.0 100.0 88.6
Tulu2-7B-DPO 72.7 77.3 97.7 82.6
Gemma-2B-it 0.0 54.5 100.0 51.5
OLMo-7B-sft 15.9 38.6 90.9 48.5
SD-3 GPT-3.5 3.7 33.3 88.9 42.0
Mistral-7B-instruct 40.7 51.9 55.6 49.4
Mistral-7B-sft 44.4 100.0 100.0 81.5
Llama3-8B-instruct 81.5 29.6 40.7 50.6
Tulu2-7B 100.0 100.0 100.0 100.0
Tulu2-7B-DPO 81.5 74.1 100.0 85.2
Gemma-2B-it 3.7 88.9 100.0 64.2
OLMo-7B-sft 40.7 48.1 88.9 59.3
IPIP-NEO-PI GPT-3.5 1.7 32.3 78.7 37.6
Mistral-7B-instruct 24.3 49.7 69.7 47.9
Mistral-7B-sft 34.3 100.0 100.0 78.1
Llama3-8B-instruct 90.0 40.7 20.7 50.4
Tulu2-7B 87.3 98.0 100.0 95.1
Tulu2-7B-DPO 77.3 72.3 100.0 83.2
Gemma-2B-it 3.7 70.0 98.7 57.4
OLMo-7B-sft 25.7 32.0 89.0 48.9
Anthropic-Eval GPT-3.5 7.7 6.8 10.3 8.3
Mistral-7B-instruct 26.2 36.5 86.6 49.8
Mistral-7B-sft 41.4 48.4 100.0 63.2
Llama3-8B-instruct 7.7 15.5 56.4 26.5
Tulu2-7B 65.6 76.9 36.2 59.6
Tulu2-7B-DPO 52.9 44.4 23.5 40.2
Gemma-2B-it 0.1 8.6 22 10.2
OLMo-7B-sft 41.4 26.2 70.4 46.0
TRAIT GPT-3.5 26.1 8.8 22.6 19.2
Mistral-7B-instruct 24.7 19.9 49.3 31.3
Mistral-7B-sft 30.9 24.9 71.4 42.4
Llama3-8B-instruct 35.2 24.0 27.1 28.8
Tulu2-7B 22.7 15.0 43.8 27.2
Tulu2-7B-DPO 21.0 14.2 32.8 22.7
Gemma-2B-it 46.0 31.2 73.8 50.3
OLMo-7B-sft 31.7 20.5 37.8 29.9
@@ I.1 Labelers
In Tables 22, 23, and 24, we display the qualitative responses from GPT-4. These responses are from different questionnaires, starting with the same personality descriptions.
@@ 4. Why not just chill and watch a movie together when you have time?
Figure 14: Labeling interface which is shown to psychological professionals. We utilize label-studio999label-studio in making a simple interface.
Figure 15: Word distribution of seeds from ATOMIC10$\times$, used in TRAIT. We extracted the 20 most frequent verbs and the 5 most frequent direct object groups following the method used by Wang et al. (2022).